OpenStudy (anonymous):

Solve for the variable in this equation. 3(x+2/5)=x-1/3

OpenStudy (anonymous):

3(x+2/5)=x-1/3 3x+6/5=x-1/3 45x+18=15x-5 30x=-23 x=-23/30

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

3(x+2/5)=x-1/3 Expand it. 3x+6/5=x-1/3 3x+6/5 -x = x- 1/3 -x 2x + 6/5 = -1/3 2x+(6/5)-(6/5)=(-1/3)-(6/5) 2x=-23/15 2x/2=(-23/15)/2 x=-23/30